# Unveiling Molecular Moieties through Hierarchical Grad-CAM Graph Explainability

### v4

@@ 1 Background
Next sections of the paper are arranged as follows: Section 2 describes the creation of the two data sets used for training the networks, their structure, and the development of the explainability algorithm. Section LABEL:sec3 reports the classification performance analysis of the implemented GNNs with respect to different metrics, and the results obtained from HGE, along with a discussion of the interpretability of these results. Finally, Section LABEL:sec5 sets the groundwork for future studies.
@@ 2.1.2 Data set 2: ChEMBL_over
Each small molecule downloaded using the API was labeled based on the threshold values given in the literature. The molecules obtained in this way have a very low active/inactive ratio. For this reason, the ChEMBL_over data set was increased using an oversampling algorithm. This is a very common practice in machine learning to compensate for unbalanced data, and it has been adopted in recent years also in the chemical domain [8, 12]. We used a Random Over Sampling algorithm in the Imbalanced Learn library [19]. Overampling was applied on inactive molecules, until an active/inactive ratio of 1:15 was obtained. The generated molecules were carefully checked and validated through the RDKit library [18] to avoid introducing a classification bias due to the presence of chemically invalid molecules. Table 1 reports the amount of inactive molecules in the ChEMBL_over data set, arranged as “Active”, “I.C.” (Inactive Compound) and “I.O.” (Inactive Oversampled).
@@ 3 Results
Experiments were arranged in three stages. At first, a GCNN classifier was trained for each target, to allow HGE computation. Furthermore, the explainability analysis was performed on 143 known kinase inhibitors from DrugBankhttps://go.drugbank.com/), and about 80% of them were correctly assigned to the annotated target. Finally, our GCNN architectures were compared with Chemprop, one of the most widely used frameworks for training message-passing neural networks aimed at predicting molecular properties.
@@ 3.2 Results on EMBER data set
The results on the EMBER data set for each single-target classifier are shown in Table 2 and Table 3. Table 2 shows the results achieved by our architecture for each target, and the last two columns show the sensitivity comparison with Chemprop. The comparison with Chemprop is reported in Table 4 where all metrics averaged on the 20 targets are reported for both approaches.
@@
Protein Bal. Accuracy F1-Score AUC G-Sensitivity C-Sensitivity
ACK 0.891 0.979 0.983 0.811 0.657
ALK 0.929 0.972 0.987 0.893 0.742
CDK1 0.921 0.956 0.969 0.907 0.650
CDK2 0.932 0.940 0.971 0.950 0.721
CDK6 0.916 0.989 0.988 0.846 0.772
CHK1 0.923 0.954 0.977 0.912 0.604
CK2A1 0.942 0.979 0.989 0.911 0.705
CLK2 0.938 0.929 0.989 0.965 0.881
DYRK1A 0.913 0.983 0.986 0.847 0.730
EGFR 0.930 0.982 0.990 0.881 0.749
ERK2 0.927 0.941 0.977 0.935 0.782
GSK3B 0.903 0.970 0.978 0.847 0.615
INSR 0.893 0.974 0.977 0.823 0.430
IRAK4 0.930 0.953 0.979 0.931 0.619
ITK 0.928 0.921 0.981 0.958 0.814
JKA2 0.945 0.945 0.984 0.960 0.786
JNK 0.918 0.923 0.971 0.946 0.691
MAP2K1 0.944 0.969 0.991 0.926 0.807
MELK 0.916 0.969 0.982 0.874 0.635
PDK1 0.913 0.973 0.986 0.860 0.693
mean 0.899 0.704
*G-Sensitivity = Our GCN Sensitivity; C-Sensitivity = Chemprop Sensitivity
@@
Table 4: Average evaluation metrics for Chemprop and our GCNN on EMBER data set
The high Sensitivity indicates a good ability of our GCNNs to correctly prioritize active ligands, and this is confirmed by the optimal performance obtained in terms of TP/P and EF values.
@@ 3.3 Results on ChEMBL_over data set
Protein Bal. accuracy AUC F1-score G-Sensitivity* C-Sensitivity*
ACK1 0.972 0.992 0.954 0.947 0.866
ALK 0.979 0.996 0.978 0.957 0.944
CDK1 0.949 0.995 0.933 0.901 0.939
CDK2 0.953 0.985 0.862 0.929 0.907
CDK6 0.952 0.951 0.901 0.914 0.893
CHK1 0.969 0.991 0.965 0.939 0.907
CK2A1 0.986 1.000 0.986 0.972 0.896
CLK2 0.956 0.991 0.954 0.911 0.899
DYRK1A 0.943 0.974 0.871 0.903 0.723
EGFR 0.914 0.961 0.715 0.887 0.895
ERK2 0.954 0.979 0.803 0.950 0.893
GSK3B 0.958 0.984 0.876 0.937 0.869
INSR 0.981 0.990 0.981 0.962 0.928
IRAK4 0.993 0.999 0.993 0.986 0.952
ITK 0.997 1.000 0.997 0.995 0.950
JAK2 0.993 0.998 0.980 0.989 0.970
JNK3 0.942 0.979 0.707 0.959 0.819
MAPK2K1 0.964 0.989 0.963 0.928 0.903
MELK 0.992 0.999 0.992 0.984 0.833
PDK1 0.988 1.000 0.988 0.977 0.947
mean 0.946 0.897
*G-Sensitivity = Our GCN Sensitivity; C-Sensitivity = Chemprop Sensitivity
@@
Table 7: Average evaluation metrics for Chemprop and our GCNN on ChEMBL_over data set.
The performance of our GCNNs on the ChEMLB_over data set remains consistent with the one obtained on the EMBER one. Our architectures performs better than Chemprop on 18 out 20 targets in terms of Sensitivity, but also in the target when they perform less than Chemprop the difference is very subtle. The TP/P and EF scores are high on all the targets. Finally, Chemprop reports an average F1-score that is higher than our architectures. This is not a limitation because our architectures were trained for reaching high Sensitivity as it is demonstrated by the average of such a metric on all the target. As a consequence, our networks perform better than Chemprop in prioritizing True Positive. This result could reveal a vulnerability in the Chemprop net, which seems more prone than our GCNNs to be misled in assessing True Positives by the presence of over-sampled molecules.
@@ 3.4 Explainability results
As shown in the table 8, only 30/44 (0.6818) drugs with multi-target activity are correctly classified, but this result can be interpreted positively. In fact, the same molecules were tested with HGE, demonstrating the robustness in correctly classifying molecular scaffolds. All the drugs have been tested with HGE, in order to assess whether the core pharmacophore could be predicted correctly, and these results are reported in supplementary material 3. Some examples are shown in figure 2, where the prediction of the molecular moieties of Apigenin, a drug that is bioactive on CDK6 and CK2A1, remains congruous, despite the bioactivity is calculated on two different targets. This result demonstrates that HGE is consistent and robust when discriminating between compounds even when the neural network is trained on different protein targets.
Figure 2: Predictions obtained with HGE for Apigenin on CDK6 and CK2A1
@@
Figure 3: Predictions obtained with HGE for Trilaciclib on CDK2 and CDK6
@@
Figure 4: Chemical features identified by HGE in a) Apigenin and Chrysin when assigned to CDK6 protein and b) Chemical features identified by HGE in Pacritinib and SB1578 when assigned to JAK2 protein
@@ 3.5 Discussion
Experimental results show clearly that HGE performs very well in real applications with respect to both the prediction task and the explanation of the molecular moieties. The main strength of the presented approach is the combination of different explanation layers (from single atoms to complex structures). Each layer is capable of catching different aspects of the molecule under investigation, and each step enriches the others, giving a complete interpretation of the molecular moieties useful to be classified as active for a specific target. As demonstrated in Figures 2, 3 and 4, the moieties are predicted consistently across similar chemotypes for the annotated target, and the HGE prediction is consistent and robust even for drugs with bioactivity on distinct targets.
@@ 4 Conclusion
Our experiments produced SOTA results with respect to classification and the moieties detection is highly precise. Such results set the framework as a very useful tool for Virtual Screening where our approach unveils potential hidden molecular moieties useful for binding a specific target. Finally, HGE is able to associate a given set of moieties with a particular target. As a consequence, very interesting insights can also be gained in the direction of drug repurposing if we use HGE to analyze a known drug that is known to own some of the moieties related to such a target. HGE performance is shown to be good and reliable both in predicting the same moieties for a target (figure 4) and in identifying the same molecular features of the same molecule on different targets (figure 2, 3). These two complementary aspects demonstrate, in our opinion, a good reliability of HGE in identifying the important chemical moieties for small molecule-target recognition. Future research will be dedicated to studying the ability of HGE to re-propose drugs and to understanding the link between classification properties and explainer prediction for assessing how to increase the amount of information that can be obtained from the explainer.

### v5

@@ 1 Background
Next sections of the paper are arranged as follows: Section 2 describes the creation of the two data sets used for training the networks, their structure, and the development of the explainability algorithm. Section 3 reports the classification performance analysis of the implemented GNNs with respect to different metrics, and the results obtained from HGE, along with a discussion of the interpretability of these results. Finally, Section 4 sets the groundwork for future studies.
@@ 2 Methods
The main objective of this work is to develop a method to identify which molecular substructures play a major role in characterizing a molecule as putatively active against a protein target. The analysis of the results provided by HGE may guide medicinal chemists in the design of novel therapeutics for a certain target, focusing the efforts on specific substructures by generating all the relevant decorations to increase the potency of a molecule or build novel putative lead compounds. For this purpose, we chose one of the major investigated protein families as drug targets [29], i.e. Kinase proteins, to build and test our HGE model. Kinases are essential regulators of numerous cellular processes, including signal transduction, cell cycle progression, and apoptosis, making them critical targets in cancer research and therapeutic development. Moreover, a massive amount of data is available from the literature, making kinases suitable for our work. However, the study described in this work was narrowed to a set of 20 kinases, while the choice of these proteins was described in [22]. These kinases are listed in Table S1 in the supplementary file S1.
2.1 Data preparation
@@ 2.1.2 Data set 2: ChEMBL_over
Each small molecule downloaded using the API was labeled based on the threshold values given in the literature. The molecules obtained in this way have a very low active/inactive ratio. For this reason, the ChEMBL_over data set was increased using an oversampling algorithm. This is a very common practice in machine learning to compensate for unbalanced data, and it has been adopted in recent years also in the chemical domain [9, 13]. We used a Random Over Sampling algorithm in the Imbalanced Learn library [20]. Overampling was applied on inactive molecules, until an active/inactive ratio of 1:15 was obtained. The generated molecules were carefully checked and validated through the RDKit library [19] to avoid introducing a classification bias due to the presence of chemically invalid molecules.
The validation has been performed using the SanitazeMol https://www.rdkit.org/docs/index.html function of Rdkit as reported in the official cookbook. Specifically, each molecule was tested with the catch_error parameter set to True in order to receive output on any errors. In this way, every single Canonical Smiles generated was tested by evaluating the parameters listed in the table 1.
Flag Description
SANITIZE_CLEANUP Cleans the structure by removing invalid bonds and normalizing the molecule.
SANITIZE_PROPERTIES Checks and assigns atomic properties such as valences and formal charges.
SANITIZE_SYMMRINGS Identifies symmetric rings to prepare for aromaticity determination.
SANITIZE_KEKULIZE Converts aromatic bonds into a Kekulé representation (alternating single and double bonds).
SANITIZE_SETAROMATICITY Determines which atoms and bonds are aromatic.
SANITIZE_FINDRADICALS Identifies and assigns unpaired electrons (radicals) to atoms.
SANITIZE_ADJUSTHS Adjusts the number of implicit hydrogens based on the valence of the atoms.
SANITIZE_CLEANUPCHIRALITY Removes or corrects invalid or ambiguous stereochemistry.
SANITIZE_FINDSTEREO Identifies stereogenic centers and configures stereochemistry (e.g., R/S, E/Z).
SANITIZE_SETCONJUGATION Determines the conjugation of bonds (e.g., conjugated $\pi$ systems).
SANITIZE_SETHYBRIDIZATION Assigns hybridization to atoms (e.g., sp, sp2, sp3).
SANITIZE_ALL Performs all the checks listed above.
Table 1: RDKit Sanitization Flags and their Descriptions
Once the validation phase had been completed, a global evaluation analysis of the new dataset was carried out. Specifically, a similarity study was performed between all the molecules generated with the oversampling procedure against the molecules downloaded from ChEMBL. This experiment allowed the construction of a similarity matrix, shown in figure 1, where it is highlighted how the inclusion of these molecules leads to no bias in terms of similarity, and on the other hand brings out a remarkable structural uniformity.
Figure 1: Heatmap of the similarity matrix obtained between the oversampled and ChEMBL molecules.
Table 2 reports the amount of inactive molecules in the ChEMBL_over data set, arranged as “Active”, “I.C.” (Inactive Compound) and “I.O.” (Inactive Oversampled).
@@ 3 Results
Experiments were arranged in three stages. At first, a GCNN classifier was trained for each target, to allow HGE computation. Furthermore, the explainability analysis was performed on 143 known kinase inhibitors from DrugBankhttps://go.drugbank.com/), and about 80% of them were correctly assigned to the annotated target. Finally, our GCNN architectures were compared with Chemprop and Attentive FP [43, 6, 42], two of the most widely used frameworks for training message-passing neural networks aimed at predicting molecular properties.
@@ 3.2 Results on EMBER data set
The results on the EMBER data set for each single-target classifier are shown in Table 3 and Table 4. Table 3 shows the results achieved by our architecture for each target, and the last three columns show the sensitivity comparison with Chemprop and Attentive FP. In addition, a comparison with Chemprop and Attentive FP is reported in Table 5 where all metrics averaged on the 20 targets are reported for both approaches.
@@
Protein Bal. Accuracy F1-Score AUC G-Sensitivity C-Sensitivity A-Sensitivity
ACK 0.891 0.979 0.983 0.811 0.657 0.0822
ALK 0.929 0.972 0.987 0.893 0.742 0.2532
CDK1 0.921 0.956 0.969 0.907 0.650 0.2650
CDK2 0.932 0.940 0.971 0.950 0.721 0.4099
CDK6 0.916 0.989 0.988 0.846 0.772 0.3158
CHK1 0.923 0.954 0.977 0.912 0.604 0.6443
CK2A1 0.942 0.979 0.989 0.911 0.705 0.3786
CLK2 0.938 0.929 0.989 0.965 0.881 0.0794
DYRK1A 0.913 0.983 0.986 0.847 0.730 0.2376
EGFR 0.930 0.982 0.990 0.881 0.749 0.8859
ERK2 0.927 0.941 0.977 0.935 0.782 0.7416
GSK3B 0.903 0.970 0.978 0.847 0.615 0.6639
INSR 0.893 0.974 0.977 0.823 0.430 0.2279
IRAK4 0.930 0.953 0.979 0.931 0.619 0.4862
ITK 0.928 0.921 0.981 0.958 0.814 0.1625
JAK2 0.945 0.945 0.984 0.960 0.786 0.9000
JNK 0.918 0.923 0.971 0.946 0.691 0.0154
MAP2K1 0.944 0.969 0.991 0.926 0.807 0.2381
MELK 0.916 0.969 0.982 0.874 0.635 0.5229
PDK1 0.913 0.973 0.986 0.860 0.693 0.3818
mean 0.899 0.704 0.3946
*G-Sensitivity = Our GCN Sensitivity; C-Sensitivity = Chemprop Sensitivity; A-Sensitiviy = Attentive FP Sensitivity.
@@
Table 5: Average evaluation metrics for Chemprop, Attentive FP and our GCNN on EMBER data set
As can be seen in table 3, our architecture outperforms in terms of sensitivity the two architectures used as a comparison on all 20 protein targets for the EMBER dataset. This result highlights the excellent capabilities of our GCN in the correct classification of small molecules and as can be seen from the table 4 it also has an very good ability to prioritise True Positives in terms of TP/P and EF. Finally, as shown in table 5 our GCN, overall performs better in terms of F1-score and AUC on the average of each target.
@@ 3.3 Results on ChEMBL_over data set
Protein Bal. accuracy AUC F1-score G-Sensitivity* C-Sensitivity* A-Sensitivity
ACK1 0.972 0.992 0.954 0.947 0.866 0.921
ALK 0.979 0.996 0.978 0.957 0.944 0.934
CDK1 0.949 0.995 0.933 0.901 0.939 0.832
CDK2 0.953 0.985 0.862 0.929 0.907 0.862
CDK6 0.952 0.951 0.901 0.914 0.893 0.814
CHK1 0.969 0.991 0.965 0.939 0.907 0.905
CK2A1 0.986 1.000 0.986 0.972 0.896 0.926
CLK2 0.956 0.991 0.954 0.911 0.899 0.861
DYRK1A 0.943 0.974 0.871 0.903 0.723 0.883
EGFR 0.914 0.961 0.715 0.887 0.895 0.870
ERK2 0.954 0.979 0.803 0.950 0.893 0.922
GSK3B 0.958 0.984 0.876 0.937 0.869 0.906
INSR 0.981 0.990 0.981 0.962 0.928 0.920
IRAK4 0.993 0.999 0.993 0.986 0.952 0.970
ITK 0.997 1.000 0.997 0.995 0.950 0.968
JAK2 0.993 0.998 0.980 0.989 0.970 0.975
JNK3 0.942 0.979 0.707 0.959 0.819 0.803
MAPK2K1 0.964 0.989 0.963 0.928 0.903 0.871
MELK 0.992 0.999 0.992 0.984 0.833 0.972
PDK1 0.988 1.000 0.988 0.977 0.947 0.954
mean 0.946 0.897 0.904
*G-Sensitivity = Our GCN Sensitivity; C-Sensitivity = Chemprop Sensitivity; A-Sensitivity = Attention FP Sensitivity.
@@
Table 8: Average evaluation metrics for Chemprop, Attentive FP and our GCNN on ChEMBL_over data set.
The GCNNs’ performance on the ChEMLB_over data set is consistent with that obtained on the EMBER one. Our architectures outperform Chemprop on 18 out of 20 targets in terms of sensitivity, but the difference between the two is quite minor. The TP/P and EF ratings are strong across all objectives. Finally, Chemprop’s average F1-score is higher than our architectures. This is not a constraint because our designs were trained to achieve high Sensitivity, as evidenced by the average of such a statistic across all targets. As a result, our networks outperform Chemprop at prioritising True Positive. This result may suggest a vulnerability in the Chemprop net, which appears more prone than our GCNNs to be misled in determining True Positives by the presence of over-sampled chemicals. Despite the different results on the two targets (CDK1 and EGFR) compared to Chemprop, in relation to the results obtained by attentive FP, our architecture discriminates better on all targets. This further provides additional proof of the robustness of the proposed GCN against two of the best-performing deep neural architectures in terms of classification.
@@ 3.4 Explainability results
As shown in the table 9, only 30/44 (0.6818) drugs with multi-target activity are correctly classified, but this result can be interpreted positively. In fact, the same molecules were tested with HGE, demonstrating the robustness in correctly classifying molecular scaffolds. Instead, the misclassification on the 14 multi-target drug molecules can be traced back to different reasons. From the analysis carried out, certainly, as shown by the heatmap in Fig 1. The molecules with which the model is trained do not present great chemical variability of the scaffolds, as is to be expected for kinase inhibitors with a similar mechanism of action and which therefore have quite similar pharmacophoric portions. For multi-target drugs, this structural similarity further complicates the possibility of recognizing motifs that can perfectly discriminate the molecules for a target. A second aspect that must certainly be underlined is that for some targets, training datasets contain fewer known molecules (see table 2), therefore the model has fewer examples on which to be trained and especially in the case of multi-target drugs shows lower performances. All the drugs have been tested with HGE, in order to assess whether the core pharmacophore could be predicted correctly, and these results are reported in supplementary material 3. Some examples are shown in figure 3, where the prediction of the molecular moieties of Apigenin, a drug that is bioactive on CDK6 and CK2A1, remains congruous, despite the bioactivity is calculated on two different targets. This result demonstrates that HGE is consistent and robust when discriminating between compounds even when the neural network is trained on different protein targets.
Figure 3: Predictions obtained with HGE for Apigenin on CDK6 and CK2A1. H-bond acceptor or donor groups are highlighted in yellow, hydrophobic groups are colored in blue, and the aromatic groups are green.
@@
Figure 4: Predictions obtained with HGE for Trilaciclib on CDK2 and CDK6. H-bond acceptor or donor groups are highlighted in yellow, hydrophobic groups are colored in blue, and the aromatic groups are green.
@@
Figure 5: Chemical features identified by HGE in a) Apigenin and Chrysin when assigned to CDK6 protein and b) Chemical features identified by HGE in Pacritinib and SB1578 when assigned to JAK2 protein. H-bond acceptor or donor groups are highlighted in yellow, hydrophobic groups are colored in blue, and the aromatic groups are green.
In order to further validate the abilities of our explainer, a comparison has been carried out with the GNNExplainer [39]. This explainer is one of the most widely used for graph explaining procedures. The experiments performed involved all drugs selected for testing using the best neural architecture for each target. GNNExplainer allows, through a training procedure, to obtain the node-feat-mask that describes the contribution of the features present in each node and an edge-mask that instead highlights the arcs relevant for obtaining the task. Using the visualise_graph method, this explainer proposes a visualisation of the result applied directly to the generalist graph. Considering the pharmaceutical domain such as the one being studied, this visualisation does not make it easy to interpret the outcome. For this reason, to facilitate comparison, the NetworkX library was used to produce a visualisation closer to the molecular structure. A comparison of the results obtained from the GNNExplainer vs. our HGE on Apigenin is shown in figure 6.
Figure 6: Results obtained with GNNExplainer compared with HGE on the Apigenin molecule.
On the left side of the image, the molecular graph built from the results obtained by GNNExplainer is shown. In order to increase readability, carbon atoms have been coloured in light blue and oxygen atoms in red; the arcs have also been coloured using a different colour scheme according to the result obtained by the explainer. Specifically, non-relevant arcs were highlighted in grey, while relevant arcs were highlighted in red. As can be seen, GNNExplainer detects as relevant the two arcs connecting C-5/C-6/O-16 and the arc connecting C-12/C-13. This result shows only the above-mentioned molecular moieties as relevant for classification, while not recognising the other molecular moieties. In contrast, HGE provides a more comprehensive description of the molecular moieties involved in Apigenin bioactivity classification, detecting both the atoms relevant and their chemical species by labelling the pharmacophoric features associated with each substructure.
@@ 3.5 Discussion
Experimental results show clearly that HGE performs very well in real applications with respect to both the prediction task and the explanation of the molecular moieties. The main strength of the presented approach is the combination of different explanation layers (from single atoms to complex structures). Each layer is capable of catching different aspects of the molecule under investigation, and each step enriches the others, giving a complete interpretation of the molecular moieties useful to be classified as active for a specific target. As demonstrated in Figures 3, 4 and 5, the moieties are predicted consistently across similar chemotypes for the annotated target, and the HGE prediction is consistent and robust even for drugs with bioactivity on distinct targets. As showed, in the Fig.2, the algorithm consistently identifies the aromatic regions and hydrogen bond acceptor features of the molecule, highlighting the same recognition patterns across the functional groups crucial for protein-ligand binding process. In Fig.3, compared to the previous case (Fig.2) where the recognition pattern identified the same moieties with the same importance, for Trilaciclib on CDK2 and CDK6 the algorithm slightly differentiated the role of the H-bond acceptor in N10 and O29. Probably, this difference is due to different potency magnitude of pharmacological effects of Trilacilib on CDK6 and CDK2. The molecule is indeed known for its selectivity in binding the ATP-binding cleft of CDK6 and the potency on CDK2 is lower and its effect could be indirect. As a matter of fact, HGE identifies chemical structures compatible with protein binding process, without considering pharmacological effect or drug potency. Figure 4 shows an example of how the algorithm is capable of retrieving and similarly interpreting very close chemotypes. In detail, Pacritinib and SB 1578 show how molecule size do not influence the algorithm capability in mantaining the functional group recognition consistency.
@@ 4 Conclusion
Our experiments produced very good results with respect to classification and the moieties detection is highly precise. Such results set the framework as a very useful tool for Virtual Screening where our approach unveils potential hidden molecular moieties useful for binding a specific target. Finally, HGE is able to associate a given set of moieties with a particular target. As a consequence, very interesting insights can also be gained in the direction of drug repurposing if we use HGE to analyze a known drug that is known to own some of the moieties related to such a target. HGE performance is shown to be good and reliable both in predicting the same moieties for a target (figure 5) and in identifying the same molecular features of the same molecule on different targets (figure 3, 4). These two complementary aspects demonstrate, in our opinion, a good reliability of HGE in identifying the important chemical moieties for small molecule-target recognition. In addition, the predictive comparison with GNNExplainer in Figure 6 shows the different ability of our explainer in labelling information in purely chemical terms. Future research will be dedicated to studying the ability of HGE to re-propose drugs and to understanding the link between classification properties and explainer prediction for assessing how to increase the amount of information that can be obtained from the explainer.